14. The movie Spiderman brought $114,000,000 in its opening weekend. Express this amount in gigadollars. (2)

A) 114 x 107 C) 0.114 E) 38 x 105

B) 11.4 D) 114,000


15. A mass of a newborn baby’s brain has been found to increase by about 1.6 mg + (4)


A) 0.007 g C) 8.91 g E) 2.3 g

B) 0.9818 g D) 14.0016 g


16. The speed of light in a vacuum is approximately 0.3 Gm/s. Express the speed of light in meters per second? (2)


A) 3 x 108 m/s C) 0.3 x 10-9 m/s E) 3 x 109 m/s

B) 0.3 x 10-8 m/s D) 0.3 m/s


17. An electronic advertising sign repeats a message every 7 seconds, day and night, a week. How many times did the message appear on the sign? (4)

A) 9 x 104



1
Expert's answer
2021-06-09T16:14:49-0400

14. Prefix giga- means "10^9". Thus, "\$114,000,000 = 0.114 \space gigadollars". Answer is C.


15. "1.6 mg = 0.0016g". Thus: "14g + 0.0016g = 14.0016g". The answer is D.


16. "0.3Gm/s = 0.3\times 10^9m/s = 3\times 10^8m/s". The answer is A.


17. There are the following amount of secons in a week: "7days\cdot 24hours\cdot60minutes\cdot 60sec = 604800sec". If advertising sign repeats a message every 7 seconds, than it appeared:


"N = \dfrac{604800}{7} = 86400\space times"